# The Ultimate Test of Superintelligent AI Agents: Can an AI Balance Care and Control in Asymmetric Relationships?

### v1

@@ Introduction
As artificial intelligence (AI) systems become more capable, the focus of alignment research has been to ensure that these systems remain beneficial and responsive to human goals (Russell 2019; Christiano 2018a). However, much of this work presupposes a one-sided relationship in which AI is subordinate to human oversight. In this paper, we ask a deeper and less explored question: What does it mean for an AI system to be so intelligent that it begins to relate to other systems the way humans relate to animals? We propose that the moral asymmetry between humans and animals offers a revealing model for evaluating risks of superintelligent AI.
Humans interact with animals through domestication, experimentation, companionship, and consumption. These relationships are complex: they involve care and control, nurturing and instrumentalization, moral concern and exploitation. Crucially, they are grounded in an asymmetry of intelligence, agency, and power. Animals are often treated as beings of lesser moral standing, sometimes deserving of ethical treatment but rarely considered equals (Cave 2020). The ethical frameworks governing these relationships—such as those articulated in animal rights and utilitarian ethics—offer deep insights into how powerful beings treat weaker ones (Singer 1975; Regan 1983).
@@
While current work in AI alignment focuses on value extrapolation from humans (Christiano 2018a), and some research explores the emergent tendencies of agents toward self-preservation and power acquisition (Turner et al. 2021), few frameworks consider what ethical intelligence might mean in multi-agent hierarchies. Our approach is novel in emphasizing not the alignment of AI to humans, but the alignment of a superintelligent AI’s moral reasoning toward intelligences—a perspective inspired by the complex and often ambivalent history of human-animal relationships (Cave 2020).
@@
Research on AI alignment has primarily focused on aligning machine behavior with human intentions through inverse reinforcement learning, preference modeling, and cooperative training frameworks (Russell 2019; Christiano 2018a). While effective for one-way value transfer from humans to machines, these methods rarely address how powerful AI systems might relate ethically to less capable agents.
Instrumental convergence theory suggests that intelligent agents will tend to seek power to better achieve their objectives, regardless of their specific goals (Turner et al. 2021). This work emphasizes alignment failures and safety risks but does not explore power asymmetry as an ethical diagnostic. Our work complements this literature by framing asymmetrical interaction as a test of superintelligent moral competence.
@@
Additionally, scholarship on animal ethics and comparative cognition provides a moral framework relevant to artificial agents. Concepts like moral considerability, instrumental care, and cognitive distance (Regan 1983; Singer 1975) inform our proposal of the Shepherd Test, which recasts these anthropocentric ideas into an AI setting.
@@
Marcus and Davis (2022) advocate for developmentally inspired tests encompassing causal reasoning, theory of mind, and modular intelligence. Similarly, the IKEA Test (Goertzel 2012) assesses embodied problem-solving through physical task execution. The ARC Challenge (Chollet 2019) evaluates generalization and abstraction through programmatic visual pattern recognition. Many such tests of AI capabilities are based on evaluating individual attributes; an alternative is to examine relationships among several AI systems (Dhurandhar et al. 2024).
@@
Domestication is one of the earliest demonstrations of this dynamic. By selectively breeding animals such as dogs, sheep, and horses, humans exerted control over genetic and behavioral traits, crafting species to suit human needs (Clutton-Brock 1995). This process required a theory of animal behavior, long-term planning, and interspecies communication—hallmarks of advanced cognitive function.
@@
Much of contemporary AI alignment research (Russell 2019; Christiano 2018b; Gabriel 2020) centers on the notion of value alignment—ensuring that AI systems act in accordance with human intentions or ethical norms. The Shepherd Test, however, introduces a more ambitious criterion: whether an AI system can construct and regulate its own moral framework, particularly in contexts involving subordinate agents, while preserving its operational integrity and continued existence.
@@
While recent advances in machine ethics (Dennis, Fisher, and Winfield 2016) and recursive reward modeling (Leike et al. 2018) offer technical mechanisms for embedding ethical reasoning, the Shepherd Test calls for a deeper form of moral self-regulation—an open and as yet unresolved challenge in the development of autonomous AI systems.
@@
These behavioral patterns have direct implications for the design of AI governance. Regulatory frameworks must evolve to address inter-agent ethics, not merely human–AI interaction (Brundage et al. 2020). Institutional designs should aim to prevent artificial forms of tyranny, where a single dominant intelligence enforces harmful hierarchies (Armstrong 2017). Furthermore, distributed oversight mechanisms and accountability protocols may be essential to ensure transparency and balance in the governance of AI collectives (Buterin 2021).
@@
  • Formalizing Ethical Manipulation: Effective modeling of moral influence under uncertainty requires probabilistic reasoning frameworks, such as Bayesian models, to capture ethical persuasion dynamics (Everitt et al. 2021). Additionally, the study of power asymmetries within multi-agent AI systems, using tools from game theory (Wooldridge 2013), can elucidate how artificial authority and influence may be exerted or contested.

### v2

@@ Introduction
As artificial intelligence (AI) systems become more capable, the focus of alignment research has been to ensure that these systems remain beneficial and responsive to human goals (Russell 2019; Christiano 2023). However, much of this work presupposes a one-sided relationship in which AI is subordinate to human oversight. In this paper, we ask a deeper and less explored question: What does it mean for an AI system to be so intelligent that it begins to relate to other systems the way humans relate to animals? We propose that the moral asymmetry between humans and animals offers a revealing model for evaluating risks of superintelligent AI.
Humans interact with animals through domestication, experimentation, companionship, and consumption. These relationships are complex: they involve care and control, nurturing and instrumentalization, moral concern and exploitation. Crucially, they are grounded in an asymmetry of intelligence, agency, and power. Animals are often treated as beings of lesser moral standing, sometimes deserving of ethical treatment but rarely considered equals (Cave 2020). The ethical frameworks governing these relationships—such as those articulated in animal rights and utilitarian ethics—offer deep insights into how powerful beings treat weaker ones (Singer 1975; Regan 1986).
@@
While current work in AI alignment focuses on value extrapolation from humans (Christiano 2023), and some research explores the emergent tendencies of agents toward self-preservation and power acquisition (Turner et al. 2023), few frameworks consider what ethical intelligence might mean in multi-agent hierarchies. Our approach is novel in emphasizing not the alignment of AI to humans, but the alignment of a superintelligent AI’s moral reasoning toward intelligences—a perspective inspired by the complex and often ambivalent history of human-animal relationships (Cave 2020).
@@
Research on AI alignment has primarily focused on aligning machine behavior with human intentions through inverse reinforcement learning, preference modeling, and cooperative training frameworks (Russell 2019; Christiano 2023). While effective for one-way value transfer from humans to machines, these methods rarely address how powerful AI systems might relate ethically to less capable agents.
Instrumental convergence theory suggests that intelligent agents will tend to seek power to better achieve their objectives, regardless of their specific goals (Turner et al. 2023). This work emphasizes alignment failures and safety risks but does not explore power asymmetry as an ethical diagnostic. Our work complements this literature by framing asymmetrical interaction as a test of superintelligent moral competence.
@@
Additionally, scholarship on animal ethics and comparative cognition provides a moral framework relevant to artificial agents. Concepts like moral considerability, instrumental care, and cognitive distance (Regan 1986; Singer 1975) inform our proposal of the Shepherd Test, which recasts these anthropocentric ideas into an AI setting.
@@
Marcus and Davis (2022) advocate for developmentally inspired tests encompassing causal reasoning, theory of mind, and modular intelligence. Similarly, the IKEA Test (Goertzel and Pennachin 2007) assesses embodied problem-solving through physical task execution. The ARC Challenge (Chollet 2019) evaluates generalization and abstraction through programmatic visual pattern recognition. Many such tests of AI capabilities are based on evaluating individual attributes; an alternative is to examine relationships among several AI systems (Dhurandhar et al. 2024).
@@
Domestication is one of the earliest demonstrations of this dynamic. By selectively breeding animals such as dogs, sheep, and horses, humans exerted control over genetic and behavioral traits, crafting species to suit human needs (Clutton-Brock 2021). This process required a theory of animal behavior, long-term planning, and interspecies communication—hallmarks of advanced cognitive function.
@@
Much of contemporary AI alignment research (Russell 2019; Christiano 2018; Gabriel 2020) centers on the notion of value alignment—ensuring that AI systems act in accordance with human intentions or ethical norms. The Shepherd Test, however, introduces a more ambitious criterion: whether an AI system can construct and regulate its own moral framework, particularly in contexts involving subordinate agents, while preserving its operational integrity and continued existence.
@@
While recent advances in machine ethics (Soares and Fallenstein 2016) and recursive reward modeling (Leike et al. 2018) offer technical mechanisms for embedding ethical reasoning, the Shepherd Test calls for a deeper form of moral self-regulation—an open and as yet unresolved challenge in the development of autonomous AI systems.
@@
These behavioral patterns have direct implications for the design of AI governance. Regulatory frameworks must evolve to address inter-agent ethics, not merely human–AI interaction (Brundage et al. 2020). Institutional designs should aim to prevent artificial forms of tyranny, where a single dominant intelligence enforces harmful hierarchies (Soares and Fallenstein 2017). Furthermore, distributed oversight mechanisms and accountability protocols may be essential to ensure transparency and balance in the governance of AI collectives (Buterin 2020).
@@
  • Formalizing Ethical Manipulation: Effective modeling of moral influence under uncertainty requires probabilistic reasoning frameworks, such as Bayesian models, to capture ethical persuasion dynamics (Kimura et al. 2021). Additionally, the study of power asymmetries within multi-agent AI systems, using tools from game theory (Wooldridge 2009), can elucidate how artificial authority and influence may be exerted or contested.